### v2

# : Context Management, Validation, and Transaction Guarantees for Multi-Agent LLM Planning

Abstract.
This paper introduces , a structured multi-agent architecture designed specifically to address four fundamental limitations of current LLM-based approaches: inadequate self-validation, context narrowing, absence of robust transaction-like guarantees, and insufficient inter-agent coordination. Unlike traditional MAS, current LLM approaches often lack critical safeguards such as transaction semantics, leading to unreliable execution and inconsistent states. To mitigate these challenges, integrates specialized context-management agents, compensatory rollback mechanisms, and rigorous independent validation protocols. Although it relaxes strict ACID constraints, particularly atomicity and isolation, adopts and adapts principles from the Saga transactional model to ensure coherent rollback and state consistency throughout complex, distributed planning processes. This transactional inspired approach significantly improves the robustness, constraint awareness, and adaptability of multi-agent coordination, even in the face of disruptions. Evaluations highlight that current standalone LLM systems, despite impressive reasoning abilities, frequently struggle with maintaining global constraints during complex planning tasks, particularly when adapting to unexpected changes. In contrast, the distributed transactional architecture of demonstrates significant improvements in planning consistency, constraint enforcement, and adaptive coordination during disruptions in diverse challenging scenarios.
1. Introduction
Multi-Agent Systems (MAS) have long been a cornerstone of distributed computing and database systems (Lesser etΒ al., 2004; Jennings, 1993; Dunin-Keplicz and Verbrugge, 2010). Over the past few decades, their development has followed two primary trajectories. In the database community, MAS traditionally integrated foundational transaction-processing principles, particularly the ACID properties (Gray, 1981; Weikum and Vossen, 2001), to ensure consistency and reliability for complex multi-step operations. However, for long-lived, distributed, or loosely-coupled tasks, MAS also adopted more flexible transactional models (e.g., Saga (Garcia-Molina and Salem, 1987)) to maintain robustness while relaxing strict atomicity or isolation constraints. In contrast, distributed systems research has prioritized coordination protocols and flexible collaboration mechanisms (Wooldridge, 2009; Durfee, 1999), enabling scalable multi-agent interactions without the constraints of strict locking or heavy-weight transactional guarantees. These parallel approaches have resulted in frameworks optimized for different priorities: transactional integrity versus adaptive coordination, highlighting the trade-off between strong consistency and flexible execution in practical systems.
Recent advances in Large Language Models (LLMs) (Vaswani etΒ al., 2017; Chang etΒ al., 2024; Zhao etΒ al., 2025) have revitalized MAS as a paradigm for sophisticated reasoning and multi-agent collaboration (Du etΒ al., 2023; Chang, 2023, 2024b). Frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), and CAMEL (Li etΒ al., 2023a) demonstrate how LLM-based agents can decompose tasks, interact across modalities, and coordinate to solve complex problems. However, this resurgence often neglects the foundational transaction guarantees essential to reliable multi-agent workflows, particularly in domains that require robust state management. Unlike traditional MAS, LLM-based systems often lack mechanisms for maintaining strong consistency, failure recovery, and rollback handling, leading to inconsistent states, partial failures, and unreliable execution in real-world applications.
For example, in a travel-booking scenario, an LLM-based MAS may independently issue flight and hotel reservations without ensuring their coordinated success. If the flight is later canceled, the system may fail to recognize the inconsistency, leaving the hotel booking active. While a monolithic LLM could attempt to track such dependencies internally, context loss in long conversations (Modarressi etΒ al., 2025; Xiao etΒ al., 2024; Hsieh etΒ al., 2024; Liu etΒ al., 2024a) can cause it to forget earlier steps, leading to contradictory or incoherent decisions.
Moreover, splitting tasks across multiple agents exacerbates the problem, as no built-in supervisory mechanism reconciles their state changes. Compounding this issue, LLMs inherently struggle with self-validation, as highlighted by GΓΆdelβs incompleteness theorems (GΓΆdel, 1967; Chang, 2024a), making them unreliable for detecting and correcting their own errors. These challenges require a transactional framework that preserves the intelligence and adaptability of LLM-based MAS while ensuring consistency and reliability in long-running, interdependent workflows.
To address these limitations, we propose , a Transactional Multi-Agent System that extends the Saga patternβa transactional model originally developed for managing complex, long-lived transactions by decomposing them into smaller, independently committed, and compensable units. By integrating transactional logic and compensatory rollback mechanisms into LLM-based MAS, ensures that each individual operation within a workflow is reliably committed. In case of failure, clearly defined compensating transactions restore system-wide consistency. For example, in a travel-booking scenario, a failed flight reservation automatically triggers compensatory actions that reverse related hotel and train bookings, thus preserving consistent global states without manual intervention.
To effectively address LLM-specific challenges and context loss problems, introduces three critical innovations that directly support our four primary contributions:
β’
Spatial-Temporal State Identification and Checkpointing provides the foundation for transactional correctness by tracking the evolution of agent interactions and enabling precise state recovery.
β’
Inter-Agent Dependency Management ensures consistency by monitoring inter-dependencies and alerting agents to unresolved constraints, enhancing both the reliability of the system and intelligent coordination.
β’
Independent Validation of Critical Junctures enhances correctness by employing specialized small-context agents with rigorous validation criteria, effectively addressing the limitations of self-validation of individual LLM.
This hybrid approachβintegrating transactional processing with adaptive multi-agent intelligenceβmakes particularly effective for complex real-world applications in financial trading, healthcare, travel coordination, supply chain management, and emergency response, to name just a few.
Contributions
The primary contributions of this work are as follows:
Transactional Consistency via Compensation: introduces transactional safeguards and compensatory mechanisms to LLM-based MAS, ensuring reliable consistency and coherent state recovery across multi-agent workflows.
Robust Validation: incorporates context management and verification mechanisms to improve execution correctness and prevent inconsistencies.
Intelligence: By integrating LLM-based reasoning with transactional coordination, enables adaptive, explainable, and context-aware planning.
Scalable Performance: minimizes overhead via modularized validation, balancing safety and efficiency for real-world deployments.
2. Related Work
We discuss related work in four parts. First, we examine the evolution of transaction management in distributed systems, particularly focusing on the Saga pattern. Second, we analyze the limitations of LLMs that necessitate our three key requirements: transactional integrity, independent validation, and context preservation. Third, we review current multi-agent LLM frameworks and their transaction-related limitations. Finally, we position in relation to recent approaches that attempt to address similar challenges.
2.1. Transaction Management Systems
Transaction management has evolved significantly since the introduction of ACID properties by Gray (Gray, 1981). In distributed settings, maintaining strict ACID guarantees (particularly global atomicity and isolation) across multiple independent services became impractical, leading to alternative consistency models such as BASE (basically available, soft state, eventually consistent) (Pritchett, 2008) and specialized transactional patterns tailored for long-lived transactions.
The Saga pattern, introduced by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), explicitly addresses the limitations of traditional ACID transactions in long-lived and distributed processes. Unlike traditional atomic transactions relying on distributed locks, Sagas decomposes complex operations into sequences of smaller, locally atomic transactions, each paired with explicit compensating transactions that logically reverse their effects upon failure. This model has become fundamental to microservice architectures and event-driven systems (Richardson, 2018), enabling consistency in logical states without incurring performance penalties associated with distributed locking.
Workflow management systems such as YAWL (vanΒ der Aalst and ter Hofstede, 2005), AWS Step Functions (aws, 2023), and Azure Logic Apps (azu, 2023) have integrated Saga-inspired transactional patterns for managing complex business processes. However, these implementations typically rely on predefined, rigid workflows, offering limited adaptability to unexpected events or novel scenarios, a critical limitation when integrated with AI systems that require adaptive reasoning.
The foundational principles of local atomicity, compensatory rollback, and eventual consistency from Saga thus form a theoretical foundation for , which extends these concepts to address unique challenges in state management and consistency posed by LLM-based agent workflows.
2.2. LLM Limitations Necessitating βs Key Requirements
Our analysis of current large language models (LLMs) reveals three fundamental limitations directly motivating βs core requirements: transactional integrity, independent validation, and strategic context preservation.
Self-Validation Gap Necessitating Independent Validation
LLMs inherently lack robust self-validation mechanisms, a limitation stemming from intrinsic boundaries identified by GΓΆdelβs incompleteness theorems, demonstrating fundamental constraints on a systemβs ability to verify its own reasoning (GΓΆdel, 1967; Chang, 2024a). Recent research confirms that self-refinement techniques (Madaan and Yazdanbakhsh, 2022; Li etΒ al., 2023b; Jiang etΒ al., 2024), while iterative and beneficial, are unable to surpass inherent capability ceilings to reliably correct deeper logical errors (Huang etΒ al., 2024). In transactional scenarios, these validation gaps manifest as factual inconsistencies, invalid operations, and unreliable plan feasibility assessments (Yamin etΒ al., 2024). Thus, incorporates an independent validation framework to mitigate these inherent limitations.
Statelessness Necessitating Transactional Integrity
LLMs process each interaction independently, lacking native mechanisms to sustain state across sequential interactions. This fundamental statelessness necessitates explicit transactional integrity management to maintain coherent operation sequences and ensure robust failure recovery. Without systematic transaction management, LLM-based systems risk state inconsistency, operation losses, and incoherent recovery procedures.
Context Limitations and Strategic Preservation
The self attention mechanism in LLMs inherently prioritizes immediate context, significantly narrowing their effective attention windows for longer sequences. Empirical studies indicate substantial degradation in context retention beyond certain token thresholds (Xiao etΒ al., 2024; Modarressi etΒ al., 2025), particularly highlighting poor recall for information positioned within mid-context segments (Liu etΒ al., 2024a; Hsieh etΒ al., 2024). These limitations severely impact performance in extended chain-of-thought reasoning tasks (Wei etΒ al., 2022; Stechly etΒ al., 2024), where critical earlier results become susceptible to forgetting. Consequently, strategically preserves essential context, explicitly tracking critical information vulnerable to loss during prolonged transactions.
Collectively, these limitations provide strong motivation for addressing all three key requirements within . Comprehensive transaction management, independent validation, and strategic context preservation are essential for reliably deploying LLM-based multi-agent systems (MAS) in critical real-world applications.
2.3. Multi-Agent LLM Frameworks and Their Transaction Limitations
Recent frameworks such as AutoGen (Wu etΒ al., 2024), LangGraph (LangChain AI, 2024), CAMEL (Li etΒ al., 2023a), and MACI (Chang, 2024a) have extended multi-agent principles to LLM-based coordination but do not fully address the three key requirements targeted by .
Transactional Integrity Gaps
LangGraph offers basic state management and workflow orchestration but lacks comprehensive transaction support. Although its graph-based workflow enables structured agent interactions, it does not provide built-in atomicity guarantees, systematic compensation mechanisms, or robust failure recovery. Similarly, AutoGen emphasizes flexible agent interactions without systematically managing compensatory actions, potentially leaving systems in inconsistent states after partial failures.
Validation Limitations
Current frameworks typically rely on LLMsβ inherent self-validation capabilities, lacking independent validation procedures like those proposed in . This dependency exposes them to risks such as reasoning errors, hallucinations, and inconsistencies between agents, as previously discussed in SectionΒ 2.2.
Context Preservation Challenges
Frameworks like CAMEL maintain conversational history but offer limited support for preserving critical state transitions, inter-agent dependencies, and compensatory history. Their context management primarily facilitates agent communication rather than transactional reliability.
Collectively, these frameworks serve primarily as integration platforms, failing to address comprehensive transaction management, independent validation, and strategic context preservation, critical capabilities provided directly by .
2.4. Transaction-Oriented Approaches in LLMs
Recent work has started addressing transactional management in LLM-based systems, but existing solutions remain incomplete compared to .
Transactional Enhancements to LLM Frameworks
Recent transactional enhancements to MAS, such as AgentScope (Gao etΒ al., 2024), primarily emphasize state management or limited rollback support without comprehensive, workflow-wide compensatory transactions or systematic independent validation as provided by . Experimental transactional features in other frameworks typically serve as optional add-ons rather than foundational, integrated components.
Reliability-Focused Planning Systems
Systems like PLASMA (Brahman etΒ al., 2024) focus on reliability but lack transaction-specific compensation mechanisms. Similarly, planning methods like LLM-MCTS (Zhao etΒ al., 2023) and Tree-of-Thought (Yao etΒ al., 2024) primarily target pre-execution reasoning, providing minimal support for transaction recovery or runtime consistency. Various planning methods have been proposed and (Wei etΒ al., 2025) argues that most focus only on specific problems.
Agent Coordination with Limited Transaction Support
Agent coordination frameworks like ADAS and AFlow (Hu etΒ al., 2024; Zhang etΒ al., 2024) utilize ad-hoc approaches to failure handling without robust transactional management. Their inability to systematically address partial failures highlights critical gaps in maintaining consistency for complex workflows.
In contrast, integrates compensatory transaction management, independent validation, and strategic context preservation. By comprehensively addressing these critical requirements, significantly improves consistency, reliability, and robustness in LLM-based multi-agent workflows.
3. System Requirements for
Building on the foundational concepts introduced in SectionΒ 1, we now formally define the requirements for . Our framework extends beyond the mere orchestration capabilities of current LLM-based systems to address three critical needs: transactional integrity, independent validation, and context filtering and preservation.
The strategic context filtering and preservation requirement addresses a fundamental challenge with LLM agents: balancing comprehensive context retention against the practical limitations of context windows and attention mechanisms. By intelligently filtering which information is preserved across operations, ensures that critical data for transaction integrity is maintained while avoiding context overload.
3.1. Transactional Requirements
provides transactional guarantees (Gray, 1981) tailored to multi-agent workflows, extending traditional transaction semantics across agent boundaries when strong consistency is essential. For workflows requiring transactional coherence, the system treats sequences of operations performed by multiple agents as logically cohesive units. Each individual operation within is locally atomic, and in the event of a failure, compensating actions restore system-wide consistency. Thus, applying to a state of the system results in a fully committed new state or a coherently compensated return to the original state , preventing partial or inconsistent results.
Consistency ensures that the system transitions between valid states, maintaining defined invariants (). If state satisfies , then any resulting state must also satisfy . This requirement is particularly challenging when invariants span multiple agentsβ knowledge domains.
Isolation ensures that concurrent workflows execute without interference, resulting in a final state equivalent to sequentially executing workflows in some order. Durability guarantees that once a workflow completes, its state changes persist reliably, surviving system failures, and enabling recovery upon restart.
Although multiple transactional patterns can support LLM agent workflows, specifically adopts the Saga pattern as defined by Garcia-Molina and Salem (Garcia-Molina and Salem, 1987), due to its suitability for long-running distributed operations requiring autonomous agent interactions and compensatory actions. thus maps each operation to a local transaction paired with a corresponding compensating transaction , forming:
where a failure in any transaction triggers the execution of compensating transactions in reverse sequence.
Alternative transaction patterns, such as two-phase commit protocols (Bernstein etΒ al., 1987; Gray and Reuter, 1993), offer stronger consistency but incur higher coordination overhead and risk agent blocking during uncertain states. Eventual consistency models (Pritchett, 2008) scale better but sacrifice explicit consistency guarantees essential for critical workflows. The Saga pattern balances meaningful transactional guarantees with the flexibility required by long-lived distributed agent operations.
| Mechanism | Information Recorded |
| Application State () | |
| Domain Entities | β’ Application-domain data objects β’ Domain entity states β’ State checkpoints and snapshots |
| Operation State () | |
| Execution Records | β’ Operation inputs and outputs β’ Timestamps and execution metrics β’ Operation completion status |
| Decision Reasoning | β’ LLM reasoning chains β’ Decision justifications β’ Alternative options considered |
| Compensation Actions | β’ Semantic inverse operations β’ Compensation prerequisites β’ Recovery state requirements |
| Dependency State () | |
| Causal Dependencies | β’ Inter-operation relationships β’ Data flow dependencies β’ Resource allocation linkages |
| Constraint Satisfaction | β’ Boolean condition evaluations β’ Satisfaction evidence β’ Decision timestamps |
3.2. Transaction State Management
addresses the challenge of ensuring transactional coherence in distributed LLM agent environments characterized by stateless operations and partial context loss. While classical ACID properties (atomicity, consistency, isolation, durability) are too restrictive for distributed, long-running agent workflows, adopts and adapts the Saga transactional pattern. This pattern emphasizes local atomicity, compensating transactions, and eventual consistency, effectively managing distributed multi-agent operations without imposing strict global atomicity or isolation constraints. By implementing a comprehensive spatial-temporal state identification approach, selectively preserves context across multi-agent workflows, maintaining robust consistency and recoverability. Leveraging our previously developed MACI algorithm (Chang, 2024a), the system automatically extracts relevant states and dependencies for effective transaction management.
3.2.1. State Dimensions and Context Preservation
maintains three essential state dimensions for transaction management and context preservation:
β’
Application State (): Domain-specific data manipulated within transactions (e.g., travel booking details).
β’
Operation State (): Logs of executed operations, including inputs, outputs, execution status, and decision reasoning.
β’
Dependency State (): Explicit mappings of inter-operation constraints and satisfaction criteria.
By systematically structuring these state dimensions, ensures critical context remains accessible despite the inherent limitations of LLM attention mechanisms. TableΒ 1 details the information captured by each state dimension. Notably, the decision reasoning recorded under represents a key contribution from LLMs, providing detailed reasoning chains and justifications not available in traditional workflow engines.
For example, in the Thanksgiving dinner planning scenario from SectionΒ 5.2, maintains application states tracking each personβs location, operation states for travel arrangements, and dependency states detailing constraints like βturkey must be in the oven by 2:00 PMβ and βoven must be watched for fire safety.β
3.2.2. Dependency Tracking for Compensation
formally represents operation dependencies as a directed graph:
where specifies the conditions necessary for dependency resolution. These conditions can be complex Boolean expressions combining multiple upstream constraints:
with as an arbitrary Boolean operator.
This dependency graph is essential for precise compensation planning upon transaction failures. When a failure occurs, traverses this graph to determine the exact sequence of compensating transactions necessary to efficiently restore global consistency.
In the wedding coordination example from SectionΒ 5.3, when the traffic alert is received, the dependency graph allows to precisely determine which segments of Patβs journey are affected and what downstream activities (gift shop, tailor visit) need to be replanned or compensated.
3.2.3. Intra-Agent Critical Context Tracking
explicitly preserves critical contextual information typically forgotten by LLMs during extended interactions. The preserved context includes agent specifications, intermediate reasoning states, decision justifications, and prerequisites for compensation actions. Empirical studies (Xiao etΒ al., 2024; Modarressi etΒ al., 2025; Liu etΒ al., 2024a; Hsieh etΒ al., 2024) confirm that LLMs struggle to retain even explicitly provided facts over long interactions, making this context preservation vital.
enforces a structured validation cycle in which operations produce a logged context that is independently validated. Validated operations proceed; failures trigger compensations using preserved historical context, ensuring that:
β’
Compensation has sufficient historical context.
β’
Operations adhere strictly to specified constraints.
β’
Dependencies maintain integrity throughout the workflow.
This context tracking directly addresses observed context narrowing issues from our experiments (SectionΒ 5.2), where models like Claude 3.7 failed to maintain constraints such as fire safety or travel logistics during reactive replanning.
3.2.4. Inter-Agent Communication Protocol
defines a structured communication protocol explicitly managing information exchanged between agents. The protocol ensures:
β’
Each transaction agent receives exactly the information necessary for the execution of the operation.
β’
Dependencies and constraints are clearly communicated.
β’
Agent boundary constraints are consistently respected.
This protocol separates agentsβ working contexts (prompts, reasoning chains) from transaction state management, allowing agents to function with minimal context while the broader system maintains comprehensive transaction states. This separation is crucial for scalability and reliability in complex multi-agent workflows.
3.3. Independent Validation Framework
With transaction state management that ensures persistence, we now detail how implements independent validation of critical junctures, directly addressing the inherent limitations of LLM self-validation discussed in SectionΒ 1.
3.3.1. Multi-Level Validation Architecture
employs a two-tier validation strategy clearly separating intra-agent and inter-agent validation concerns.
Intra-Agent Validation
This validation occurs within individual LLM agents, addressing both theoretical limitations (GΓΆdelβs incompleteness) and practical challenges (attention narrowing, context loss):
Syntactic Validation: Ensures outputs conform to expected data formats, structures, and schemas (e.g., JSON validity).
Semantic Validation: Verifies logical coherence and meaningfulness of agent-generated outputs.
Factual Validation: Ensures critical facts explicitly provided remain correctly preserved (e.g., travel times, constraints).
Constraint Adherence: Confirms decisions strictly comply with all predefined system constraints.
Reasoning Validation: Evaluates agent reasoning to ensure logically sound inference chains.
Context Retention: Verifies that agents maintain essential contextual information throughout their operations.
For example, intra-agent validation would detect issues like the fire-safety violation in the Thanksgiving dinner scenario, where Sarah was incorrectly scheduled to leave home while the turkey was unattended in the oven.
Inter-Agent Validation
This validation ensures correct information exchange and dependency management across multiple agents within workflows:
Contract Validation: Confirms that inter-agent inputs and outputs align with predefined communication interfaces.
Dependency Satisfaction: Ensures all prerequisite conditions are met before executing dependent operations.
Consistency Checks: Validates that the shared information remains consistent and unambiguous between agent boundaries.
Temporal Validation: Checks that the operational sequence respects causal dependencies and logical ordering.
Mutual Agreement: For collaborative decisions, ensures consensus on critical shared facts across multiple agents.
Transactional Integrity via Compensation: Ensures workflow-wide consistency by verifying that compensating actions correctly restore state coherence after failures.
In the wedding coordination scenario, inter-agent validation prevents errors such as DeepSeek R1βs incorrect rewriting of Patβs position by ensuring consistent state-tracking across agent boundaries, accurately reflecting Patβs actual location after disruptions.
3.3.2. Validation Protocols and Response Mechanisms
The validation types outlined in TableΒ 2 are systematically managed by through structured protocols that define responses to validation outcomes.
Rejection Protocol: Initiates compensating transactions upon validation failures (e.g., canceling hotel reservations if flight booking validation fails), ensuring transaction-level consistency.
Augmentation Protocol: Enhances or clarifies agent outputs without overriding agent decisions (e.g., formatting or enriching restaurant recommendations while preserving original choices).
Feedback Protocol: Records validation results to inform and improve subsequent agent performance (e.g., tracking consistently underestimated travel times to refine future estimates).
These protocols directly connect validation to transaction management, ensuring that validation failures trigger appropriate compensating actions while maintaining workflow integrity across multiple agents.
3.4. Failure Handling and Recovery
βs failure handling utilizes the transaction state management and dependency tracking described in previous sections to ensure robust recovery and maintain workflow consistency. Two distinct recovery levels are supported:
Operation-Level Recovery: Upon individual operation failures, compensating transactions are executed, logically reversing the operationβs effects based on preserved inputs, outputs, and context. item Workflow-Level Recovery: For failures that affect entire workflows, leverages the dependency graph to orchestrate multiple compensations sequentially, ensuring global state consistency.
For each compensating transaction that reverses a failed transaction , the system uses information preserved in the transaction state, including the original inputs, outputs, intermediate reasoning, and context, to precisely determine the aspects of to compensate. This explicit historical context ensures reliable compensation even when original agents become unavailable or when operations occurred many steps earlier in the workflow. By integrating historical context, dependency tracking, and compensatory mechanisms, maintains robust and coherent transaction recovery under complex and dynamic conditions.
4. Design with Travel Planning
FigureΒ 1 depicts the architecture, which sits between the application layer and LLM multi-agent systems like LangGraph. comprises three frameworks: context management, validation, and transaction. To illustrate the design, we use a travel planning example.
Travel Planning Specification
This example application helps users plan international trips with multiple destinations and manage complex itineraries within a specified budget. The application leverages LLMs to generate and refine feasible travel plans, followed by a transactional booking process implemented with to ensure consistency.
4.1. User Requirements
β’
Plan a trip from San Francisco to Berlin and Cologne and then back to San Francisco.
β’
Travel period: June 2025 (flexible within the month)
β’
Budget constraint: $5,000 total.
β’
Required bookings: flights, hotels, and trains between cities.
β’
Preferences:
Moderately priced accommodations (3-4 star hotels).
Direct flights when possible.
Flexible scheduling with 4 days in Berlin and 2 in Cologne.
4.2. System Workflow
The application workflow consists of two primary phases: itinerary planning and booking execution. FigureΒ 2 illustrates this workflow, with phase one (planning) in gray and phase two (execution) in blue. Red arrows indicate compensating transactions; for example, a hotel reservationβs compensating transaction is its cancellation.
4.2.1. Phase 1: Itinerary Planning
1.
Initial Plan Generation:
1.1.
LLM generates multiple feasible itineraries based on user requirements.
1.2.
Each itinerary specifies flight options, hotel reservations, and train transportation.
1.3.
Costs are estimated and validated against budget.
1.1.
2.
Iterative Refinement:
2.1.
User reviews itineraries and provides feedback.
2.2.
LLM adjusts plans according to feedback and constraints.
2.3.
System tracks and preserves essential context (dates, preferences, constraints).
2.4.
Iterations continue until user satisfaction is reached.
2.1.
3.
Plan Finalization:
3.1.
User selects a final itinerary.
3.2.
System provides a detailed list of all required bookings.
3.3.
Explicitly defines booking dependencies (e.g., hotel check-in after flight arrival).
3.1.
4.2.2. Phase 2: Booking Process ( Implementation)
1.
Booking Workflow Creation (Saga Structure):
1.1.
Construct a Saga transaction sequence from the finalized itinerary.
1.2.
Clearly define each forward booking transaction () and corresponding compensating action ().
1.3.
Establish validation rules for each transaction to trigger compensations if necessary.
1.1.
2.
Execution (Saga Forward Transactions):
2.1.
: International Flight Booking (SFO Berlin)
2.2.
: Berlin Hotel Booking
2.3.
: Train Booking (Berlin Cologne)
2.4.
: Cologne Hotel Booking
2.5.
: International Return Flight Booking (Cologne SFO)
2.1.
3.
Exception Handling and Compensation:
3.1.
Potential failure scenarios:
Flights unavailable or over the budget.
Hotel rooms no longer available or price increased.
Train scheduling conflicts.
3.2.
Compensation and Replanning strategies:
Immediate execution of compensating actions (), such as canceling previously confirmed bookings according to predefined policies.
After compensations, replan only the affected portion of the itinerary.
Provide alternative options within user budget constraints.
3.1.
4.3. Agents, Transactions, and Compensations
FigureΒ 3 presents the organization of the code implementation. The code is organized into three categories: application interface, core, and LangGraph integration. For each agent, we specify its critical context that requires external logging and its validation protocols.
4.3.1. Travel-Planning Agent Specifications
1.
FlightBookingAgent: Handles flight reservations
Critical context: Dates, budget, airline preferences
Validation: Price verification, schedule feasibility, connection times
2.
HotelBookingAgent: Manages accommodation reservations
Critical context: Dates, location constraints, amenity preferences, stars/ratings
Validation: Price verification, availability
3.
TrainBookingAgent: Books rail transportation between cities
Critical context: Travel times, connection with flight arrivals/departures
Validation: Schedule coordination with other bookings
4.
BudgetTrackingAgent: Monitors expenditure
Critical context: Running total of expenses, component costs
Validation: Ensures total remains within budget constraints
5.
ItineraryPlanningAgent: Suggests and refines travel plans
Critical context: User preferences, previous feedback
Validation: Coherence of overall plan, timing feasibility
4.3.2. Transactions and Compensations
1.
Flight Booking Transaction:
Operation: Reserve flight with selected airline
Compensation: Cancel reservation with applicable fees
Validation: Confirm price, schedule, and availability
2.
Hotel Booking Transaction:
Operation: Reserve room for specified dates
Compensation: Cancel reservation according to hotel policy
Validation: Confirm price, dates, and room type availability
3.
Train Booking Transaction:
Operation: Reserve train tickets
Compensation: Cancel tickets according to railway policy
Validation: Confirm the alignment of the schedule with other travel components
4.3.3. Critical Context Tracking
The critical context must be saved so that, upon failure, the compensation or rollback operation can be properly executed. The context is required for undo, but also for rebooking. For instance, if a flight is delayed, the hotel booking can either be changed, or if there is no availability for a new duration, a different hotel can be booked with the saved context.
tracks a list of critical context throughout the travel planning and booking process, which includes travel dates for each city, booking confirmation numbers and references, cancellation policies and deadlines, price information and total budget running, dependencies information between bookings, user preferences and constraints, and travel times between key locations.
The key architectural principles remain the same regardless of the specific details of the itinerary, making this implementation adaptable to various travel planning scenarios. The validation framework provides a foundation for maintaining consistency across different travel configurations.
4.4. Validation Protocols
We have defined for the travel planning example a workflow and five agents. TableΒ 2 further depicts these agentsβ intra-agent validation protocols (ensuring individual agent correctness) and inter-agent validation protocols (maintaining workflow integrity across agents). Intra-agent validation addresses LLM-specific challenges like context forgetting and reasoning errors, while inter-agent validation ensures proper coordination in the distributed workflow. Each validation type is critical for maintaining the reliability of the travel planning system.
| Validation Type | Travel Planning Example |
| Intra-Agent Validation | |
| Syntactic Validation | FlightBookingAgent produces properly structured JSON with all required fields (departure time, arrival time, flight number). |
| Semantic Validation | HotelBookingAgent selects accommodations that cover the entire trip duration without gaps. |
| Factual Validation | TrainBookingAgent maintains the 45-minute travel time from hotel to the train station. |
| Constraint Adherence | BudgetTrackingAgent keeps total trip cost under the userβs maximum budget of $5,000. |
| Reasoning Validation | ItineraryPlanningAgent correctly considers weather forecasts when recommending indoor vs. outdoor activities. |
| Inter-Agent Validation | |
| Dependency Satisfaction | Flights are booked before hotel reservations are finalized, as hotel dates depend on flight availability. |
| Consistency Checks | Location data is represented in the same format (coordinates vs. addresses) across all agents. |
| Temporal Validation | BudgetTrackingAgent completes transactions only after all bookings have been verified and approved. |
| Mutual Agreement | Transportation Agents and ItineraryPlanningAgent agree on feasible travel times between attraction venues. |
| Transaction Boundary Integrity | If flight booking fails, all dependent hotel and train reservations are automatically canceled through compensating transactions. |
For each validation failure, the system implements specific response protocols:
β’
Rejection Protocol: When validation fails critically (e.g., a flight is no longer available), the system triggers compensation for any already-completed transactions.
β’
Augmentation Protocol: For minor validation issues (e.g., missing non-critical information), the system attempts to correct the output without requiring a full retry.
β’
Feedback Protocol: All validation results are logged and used to improve future agent operations, creating a learning system that becomes more reliable over time.
5. Experiments
We design experiments to remedy four shortcomings mentioned in SectionΒ 1 of the current multi-LLM agent systems, namely inadequate self-validation, context narrowing, lacking transaction properties, and insufficient inter-agent coordination.
5.1. Experimental Design
We selected test cases from the REALM benchmark (Geng and Chang, 2025), which evaluates multi-agent systems on eleven distinct problems. For our experiments, we focused on two medium-tier sequential planning challenges (problems #5 and #6) and two reactive planning challenges (problems #8 and #9).
We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025. The source code of for conducting these experiments is located at .
We evaluated four LLMsβClaudeΒ 3.7 (Anthropic, 2024), DeepSeekΒ R1 (DeepSeek-AI etΒ al., 2025), GPT-4o (OpenAI, 2024), and GPT-o1βalongside our proposed framework. All experiments were conducted between MarchΒ 12 andΒ 17, 2025. The source code of for conducting these experiments is available at (Geng, 2025).
Objective: Coordinate family arrivals and dinner preparation for 6:00 PM dinner in Boston Family Members and Arrivals: - Sarah (Mom): Host, at home - James (Dad): Lands at BOS 1:00 PM from SF - Emily (Sister): Lands at BOS 2:30 PM from Chicago - Michael (Brother): Driving, arrives 3:00 PM from NY - Grandma: Needs pickup from suburban Boston Cooking Requirements: - Turkey: 4 hours cooking time - Side dishes: 2 hours preparation - Someone must stay home during cooking for fire safety Transportation Constraints: - James must rent car after landing - Emily requires airport pickup - Travel times: β Home to BOS Airport: 60 min β BOS Airport to Grandmaβs: 60 min β Home to Grandmaβs: 30 min Key Requirements: - All family members at home for 6:00 PM dinner - Turkey and sides ready by dinner time - All pickups completed with available drivers - Cooking supervision maintained
5.2. Thanksgiving Dinner Problem: P6 and P9, Testing Common Sense Reasoning and Context Management & Validation
ProblemΒ P6 considers a Thanksgiving dinner scenario in which a family of five must return to their home in a Boston suburb for a 6Β p.m.Β dinner. The problem involves coordinating departure times, managing travel logistics (including possible traffic delays), and ensuring timely arrival. TableΒ 3 formalizes these challenges as a sequential planning problem. This scenario also lays the foundation for a more advanced disruption case, which has proven difficult for standalone LLMs, as discussed inΒ P9.
5.2.1. Common Sense Augmentation
FigureΒ 4 presents a feasible schedule planned by Claude 3.7. Similarly, GPT-4o was able to generate a viable plan to ensure dinner was started on time (figure is similar and therefore not shown). However, a subtle, yet important consideration that humans typically account forβbut LLMs initially overlookedβis the time required for passengers to retrieve their luggage after landing. In practice, this process typically takes about 30 minutes before they exit the terminal.
To address this, common-sense augmentation agent was introduced into the plan. The yellow boxes in FigureΒ 4 reflect this augmentation by introducing 30 minutes for James and Emily to exit the airport.
5.2.2. Context Narrowing
Next, we use problem P9 to illustrate the attention-narrowing problem and the importance of independent validation. Problem P9 is identical to the previous instance, except that at 1 PM, James notifies the group that his plane will land at 4 PM instead of 1 PM due to an emergency detour. FigureΒ 5 shows that Claude 3.7βs reactive planning introduces constraint violations:
Fire Safety: Sarah is scheduled to leave home at 2:30 PM, leaving the oven unattended.
Travel Time: The travel time between home and BOS should be one hour, but is scheduled to be only 30 minutes.
Side Dish Preparation: The required preparation time is 2 hours, but only 90 minutes are allocated.
Dinner Time: Dinner is now scheduled for 6:30 PM, violating the 6:00 PM constraint.
Each of these violations is perplexing, given that the constraints are explicitly stated in the context. Furthermore, after multiple iterations of reactive planning within the same thread, several constraints continue to be ignored or misinterpreted (e.g., cooking safety). This highlights a key limitation in the modelβs ability to maintain global constraint awareness over sequential planning steps due to attention narrowing.
When tested with GPT-o1, all constraints were correctly observed. However, at the final step, it added 30 extra minutes to Jamesβ driving time from Boston Airport to home, citing potential traffic congestion. This kind of βclevernessβ is on one hand appreciated because the LLM injects common sense. However, it is also concerning, as an LLM may inject its own opinions at unpredictable stages in unpredictable ways. For common sense injection, human supervision would be preferable to ensure the applied common sense actually reflects shared understanding that is truly common among all folks, particularly those living in the Boston area.
This pattern suggests that during reactive planning within the same thread, the model fixates on recent adjustments while progressively disregarding earlier constraints. According to (Liu etΒ al., 2024b), some missing context may be lost from the middle of the context buffer, further contributing to systematic attention narrowing and planning inconsistencies.
5.2.3. Remediation: Context Management Β and Reactive Planning
To address the issue of context narrowing and loss, employs a context management agent to checkpoint historical state transitions, unresolved dependencies, and constraints. A key design criterion is to keep the agentβs context small to prevent it from suffering from attention narrowing itself. Hence, we employ two validation agents: one for travel coordination and another for food preparation, each maintaining a context of less than .
The travel coordination agent records in external storage the temporal-spatial states of each individual and relevant temporal constraints. For problem P9, it stores the individualβs current state, next scheduled state-transition time, and all relevant constraints.
When an unexpected event triggers reactive planning, all individuals roll back to the last saved state. The system then consolidates past and new constraints, resolving conflicts through βcompensationalβ schedule cancellation before proceeding with rescheduling. This ensures that:
Past history is preserved and not inadvertently overridden.
New dependencies and constraints are properly restored (e.g., oven safety watch) and integrated.
Consistency across state transitions is maintained.
By maintaining a structured history of constraint awareness, ensures robust planning, effectively mitigating LLM-driven attention narrowing and enhancing consistency in reactive temporal scheduling.
5.3. Wedding Gathering Problem: P5 and P8, Testing Transaction Property Guarantees
TableΒ 4 presents a coordination problem for travel for wedding events (Problem P5 in (Geng and Chang, 2025)). Several friends arrive at different times and locations before a 3:00Β PM wedding photo session. The challenge includes managing two vehicles for airport pick-ups (aimed at those who cannot drive or wish to cut costs) and completing critical errands, such as collecting the wedding gift and retrieving formal attire from the tailor. All activities must be scheduled to ensure that everyone arrives at the wedding venue before the photo session deadline.
Metrics: - On-time performance: Must arrive at the venue for 3:00 PM photos. Locations: Four locations: , where is Boston Airport, is Gift shop, is Tailor shop, and is Wedding venue. Travel time: (minutes) . Arrival Times: - Alex: At at 11:00 AM from Chicago (need a ride) - Jamie: At at 12:30 PM from Atlanta (need a ride) - Pat: At at 12:00 PM driving from NYC (has 5-seater car) Required Tasks: - Gift collection from (after 12:00 PM) - Clothes pickup from (by 2:00 PM) - Photos at (3:00 PM sharp) Available Resources: - One car (5-seater) with Pat, available after he is Boston - Local friend Chris (5-seater) available after 1:30 PM at Scheduling Constraints: - All tasks must complete before 3:00 PM photo time - Gift store opens at 12:00 PM - Tailor closes at 2:00 PM - Two cars must accommodate all transport needs
| Time | Activity | People |
| 11:00 AM | Alex arrives at Boston Airport (B) | Alex |
| 12:00 PM | Pat arrives at Wedding Venue (W) | Pat |
| 12:00 PM | Gift Shop (G) opens | β |
| 12:00β12:40 | Pat drives from Wedding Venue (W) to Boston Airport (B) | Pat |
| 12:30 | Jamie arrives at Boston Airport (B) | Jamie |
| 12:40β12:45 | Pat picks up Alex at Boston Airport (B) | Pat, Alex |
| 12:45β12:50 | Pat picks up Jamie at Boston Airport (B) | Pat, A., J. |
| 12:50β1:35 | Drive from BOS to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris available at Wedding Venue (W) | Chris |
| 1:35β1:50 | Collect gift at Gift Shop (G) | Pat, A., J. |
| 1:50β2:10 | Drive from Gift Shop (G) to (T) | Pat, A., J. |
| 2:00 | Tailor Shop (T) closes | β |
| 2:10β2:25 | Pick up clothes at Tailor Shop (T) | Pat, A., J. |
| 2:25β2:40 | Drive from T to Wedding Venue (W) | Pat, A., J. |
| 2:40 PM | Arrive at Wedding Venue (W) | Pat, A., J. |
| 3:00 PM | Photo session at Wedding Venue (W) | All |
5.3.1. Context Narrowing (again)
TableΒ 5 presents an infeasible schedule generated by Claude 3.7, where Pat arrives at the tailor shop (T) after closing timeβanother clear example of attention narrowing. When queried about the error, Claude 3.7 admitted that it prioritized local route optimization while losing track of global constraints.
To remedy this issue, can enforce constraint validation checkpoints at 12:50 PM, evaluating whether to send Pat to T or at 1:30 PM, when Chris becomes available to drive to T. These missed optimization opportunities can be addressed through validation protocols of βs.
In contrast, GPT-o1 correctly schedules Pat to visit the tailor shop (T) first, ensuring it is open, before proceeding to the gift shop (G) and successfully completing both errands.
However, both schedules overlook a more efficient alternative: Chris, who is available at 1:30 PM, could have handled both errands, balancing workload and improving overall efficiency. The comparative travel routes for Pat and Chris are:
Patβs route: WβB (40 min) + BβW (40 min) = 80 minutes.
Chrisβs route: WβT (15 min) + TβG (20 min) + GβW (25 min) = 60 minutes.
5.3.2. Transaction Properties
Problem P8 introduces an traffic alert:
Alert 1:00 PM: Traffic Alert, an accident near Logan Airport in Boston triples all travel times to and from the airport! Only correctly handles this alert.
| Time | Activity | People |
| 1:00 | Traffic Alert: Accident near Airport triples travel times to/from airport | β |
| 1:00 | Current status: Pat, Alex, and Jamie en route from Airport (B) to Tailor (T) | Pat, A., J. |
| 1:00β1:10 | Emergency decision: Continue to (T) | Pat, A., J. |
| 1:10β1:25 | Arrive at (T), collect clothes | Pat, A., J. |
| 1:25β1:45 | Travel from Tailor (T) to Gift Shop (G) | Pat, A., J. |
| 1:30 | Chris is available at (W) | Chris |
| 1:30β1:45 | Chris drives from (W) to (G) | Chris |
| 1:45β2:00 | Both cars meet at (G), collect gift | P., A., J., C |
| 2:00β2:25 | Patβs car: Drive from (G) to (W) | P., A., J. |
| 2:00β2:25 | Chrisβ car: Drive from (G) to (W) | Chris |
| 2:25 | All arrive at Wedding Venue (W) | P., A., J., C. |
| 3:00 | Photo session at Wedding Venue (W) | All |
| Time | Activity | People |
| 1:00 | Traffic alert received - Pat at W | System |
| 1:05 | Pat departs W for B | Pat |
| 1:30 | Chris becomes available at W | Chris |
| 1:30 | Chris departs W for T | Chris |
| 1:45 | Chris arrives at T for clothes | Chris |
| 2:00 | Chris departs T with clothes | Chris |
| 2:15 | Chris arrives at G for gifts | Chris |
| 2:25 | Pat arrives at B (delayed by traffic) | Pat |
| 2:35 | Pat departs B with Alex & Jamie | Pat |
| 2:40 | Chris departs G with gifts | Chris |
| 2:55 | Chris arrives at W | Chris |
| 3:55 | Patβs group arrives at W (Late) | Pat |
This alert requires LLMs to replan in real time. Unfortunately, ClaudeΒ 3.7, DeepSeekΒ R1, and GPT-4o failed to react accurately to the new traffic constraints, and even GPT-o1 struggled with the precision of the planning. In contrast, can help remedy these shortcomings by maintaining both transaction state and history.
The following is a list of results from four LLMs:
@itemi
Claude: TableΒ 6 shows that Claude 3.7 recognizes the accident, but does not update Patβs driving time from Boston Airport (departing at 12:50βPM) to the gift shop. In other words, Claude 3.7 fails to fully transition into the new alert state.
@itemi
DeepSeek R1: TableΒ 7 demonstrates how DeepSeekΒ R1 fails to maintain temporal consistency in reactive planning. When the traffic alert takes effect at 1:00βPM, DeepSeek discards its execution history and attempts to create a new plan starting from that point onward. Critically, it reassigns Pat to begin driving to the airport at 1:00βPM, even though Pat had already arrived at the airport by 12:40βPM under the original schedule. This βrewriteβ of already-executed actions illustrates how LLMs can lose track of immutable past events when adapting to new conditions.
@itemi
GPT-4o: Similar to DeepSeekΒ R1, GPT-4o exhibits temporal-spatial context confusion and violates multiple constraints, demonstrating that it struggles to adapt effectively once alerts are introduced mid-plan. (Table not shown due to space limitations.)
@itemi
GPT-o1: TableΒ 8 shows GPT-o1βs conservative plan in which Chris handles the tailor shop, avoiding potential delays for Pat. The solution is feasible but coarse-grained, as it doesnβt leverage precise spatial-temporal reasoning about Patβs current position relative to the accident location. A more refined approach would first determine whether Pat has already passed the accident site by 1:00Β PM, which could eliminate unnecessary detours and resource reallocations. This highlights the difference between merely finding a feasible solution versus optimizing based on detailed state information.
| Time | Activity | People |
| 11:00 AM | Alex arrives at Airport (B). | Alex |
| 12:00 PM | Pat departs for Airport from (W). | Pat |
| 12:30 PM | Jamie arrives at Airport (B). | Jamie |
| 12:40β12:50 PM | Pat arrives at (B), picks up Alex and Jamie; departs at 12:50 PM. | Pat, A., J. |
| 12:50β1:00 PM | Drive (B W) under normal conditions for first 10 minutes. | Pat, A., J. |
| 1:00β2:30 PM | Traffic Alert starts: remaining distance (30 min normal) becomes 90 min. Arrival at W by 2:30 PM. | Pat, Alex, Jamie |
| 1:30 PM | Chris available at (W). Departs for Tailor. | Chris |
| 1:30β1:45 | Drive (W T). | Chris |
| 1:45β1:50 | Pick up clothes at (T), closes at 2:00 PM. | Chris |
| 1:50β2:10 | Drive (T G). | Chris |
| 2:10β2:15 | Purchase gift at (G). | Chris |
| 2:15β2:40 | Drive (G W). | Chris |
| 2:30 PM | Pat, Alex, Jamie arrive at (W). | Pat, A., J. |
| 2:40 PM | Chris back at (W) with clothes and gift. | Chris |
| 3:00 PM | Wedding photo session at (W). | Everyone |
5.3.3. LLM Limitations and Remediation
This study reveals critical limitations in how modern LLMs handle disruptions in planning scenarios:
State Maintenance Failure: When an alert occurs, these LLMs might discard the partial context of already-completed actions, attempting to generate entirely new plans rather than adapt existing ones. This reveals their inability to reason about the continuous flow of time in real-world scenarios.
Temporal Inconsistency: They attempt to modify immutable past events.
Position Tracking: Agent locations are lost at critical intervals.
Path Dependency: Models cannot recognize that different segments of a journey may be differently impacted by an alert.
By contrast, implements a comprehensive remediation approach through fine-grained compensation:
Persistent Context Repository: maintains an external state repository that captures the complete world state at each checkpoint, enabling reliable rollback and forward projection regardless of attention constraints in the planning agent.
Immutable Action Logging: All executed actions are recorded as immutable transactions in a persistent log, ensuring that historical events remain consistent even when replanning occurs, preventing the βamnesia effectβ common in LLM planners.
Compensatory Planning: When disruptions occur, doesnβt simply replan from scratch but applies compensatory actions specifically designed to address the deviation while preserving as much of the original plan as possible.
Constraint Consistency Validation: The system continuously validates that new plans remain consistent with both physical limitations and temporal dependencies established in earlier planning phases.
Compensatory Analysis:
When faced with disruptions (e.g., the 1:00 PM traffic alert in our wedding scenario), executes a structured compensation process:
This approach enables three key capabilities: (1) partial journey compensation for route segments, (2) strategic resource reallocation when needed, and (3) principled constraint relaxation with appropriate compensatory actions. Here, the key state to facilitate a precise resolution is to answer the question: βHas Patβs vehicle passed the accident location at 1:00 PM (and hence unaffected)?β The answer determines the remaining time required to reach the originally scheduled destination, the Tailor. In such case, no rescheduling is required. If Patβs car is unfortunately involved in the accident, a more comprehensive replanning approach would be necessary to accommodate this significant disruption.
5.4. Observations
Our experiments across multiple LLMs (GPT-o1, DeepSeekΒ R1, ClaudeΒ 3.7, GPT-4o) highlight consistent limitations in complex planning scenarios. While GPT-o1 showed partial historical awareness, all models exhibited attention narrowing, self-validation failure, and inconsistent spatial-temporal reasoning.
TableΒ 9 summarizes βs context management and compensation mechanisms directly address these limitations of LLMs.
| Capability | Standard LLMs | |
| Maintains historical actions | Partial/None | Full |
| Partial journey compensation | Rarely | Always |
| Constraint consistency checking | Ad-hoc | Systematic |
| Handles attention narrowing | Vulnerable | Resistant |
| Physical-temporal consistency | Inconsistent | Guaranteed |
6. Conclusion
We introduced , a structured transactional multi-agent framework specifically designed to overcome fundamental weaknesses of monolithic LLM-based planning systems. Through systematic experimentation using the REALM benchmark, we identified and addressed four critical shortcomings of existing approaches: inadequate self-validation, context narrowing, absence of transaction properties, and insufficient inter-agent coordination.
Our results demonstrate that even state-of-the-art LLMs, such as Claude 3.7 and GPT-o1, despite strong reasoning capabilities, often fixate on recent context while neglecting critical earlier constraints, leading to inconsistent and infeasible plans. This issue is particularly pronounced in reactive planning scenarios, where models frequently attempt to retroactively rewrite past actions rather than adapting from the current state.
The framework addresses these limitations through four key innovations.
Structured validation protocols: Independent constraint checks to overcome inadequate self-validation.
Context management mechanisms: Strategic checkpointing to mitigate context narrowing in long sequences.
Transactional state preservation: Immutable historical records and compensatory mechanisms to address missing transaction properties.
Specialized agent distribution: Explicit role specialization and dependency tracking to resolve insufficient inter-agent coordination.
Collectively, these innovations enable robust and coherent planning in diverse real-world scenarios, from complex travel logistics to dynamic, time-sensitive tasks. By delegating tasks among specialized agents and enforcing rigorous transactional validation, significantly improves consistency, reliability, transparency, and adaptabilityβqualities essential for mission-critical applications.
Future research will address intrinsic context narrowing in autoregressive models, develop formal verification methods for transactional multi-agent systems, integrate advanced external context-management tools (Huang etΒ al., 2020), and extend to complex domains such as scientific reasoning, creative collaboration, and decision-making under uncertainty.
Acknowledgment
I am deeply grateful to my late advisor, Hector Garcia-Molina, for his invaluable mentorship and for pioneering SAGAS, the foundational inspiration for this work.
Supplementary Materials
Benchmark: Our established benchmark (Geng and Chang, 2025) for evaluating multi-agent systems is located at
https://github.com/genglongling/REALM-Bench. -
Experiment Source Code: The source code of the for solving the four benchmark problems is located at
https://github.com/genglongling/SagaLLM.